Express 1 + 2i in polar form.
Answer by venugopalramana(3286)

Z=1+2i...DIVIDE AND MULTIPLY BY |Z|
|Z|=SQRT(1^2+2^2)=SQRT(5)=R SAY.....
Z= SQRT(5)*(1/SQRT(5)+i*(2/SQRT(5))...LET
COS(A)=1/SQRT(5)...THEN......SIN(A)=2/SQRT(5)...SO
Z=R{COS(A)+iSIN(A)}=R*E^(iA)WHERE R AND A ARE AS GIVEN ABOVE ...IN POLAR FORM THEY ARE................{SQRT(5),COS INVERSE(1/SQRT(5)}